Fe(s)+O2(g)+H2O(l)=Fe(OH)3(s)

A) Balance the equation.

B) A 27.9-gram iron nail falls into a stream and remains at the bottom rusting for many years. Eventually all the iron reacts to form rust. How many grams of rust form?
Answer #1

4Fe + 3O2 + 6H2O = 4Fe(OH)3

Moles of Fe = mass / molar mass

= 27.9 / 55.8

= 0.5 moles

Now , moles of Fe = moles of Fe(OH)3

= 0.5 mole

Mass of Fe = mole * molar mass

= 0.5 * 106.87

= 53.435 grams

If you have any query please comment

If you satisfied with the solution please like it thankxx